OpenStudy (anonymous):

By what number must you multiply each side of the equation 3/4x = 9 to produce the equivalent equation x = 12? a.) 1/4 b.) -3/4 c.) 4 d.) 4/3

OpenStudy (zzr0ck3r):

hint: \(\frac{a}{b}x=c \) implies \(x = \frac{b}{a}c\)

OpenStudy (anonymous):

how would i set it up? ive tried with my dad and we still cant get it...

OpenStudy (zzr0ck3r):

\(\frac{3}{4}x=9\rightarrow \frac{4}{3}\frac{3}{4}x=\frac{4}{3}*9\rightarrow \frac{\cancel{4}}{\cancel{3}}\frac{\cancel{3}}{\cancel{4}}x=\frac{4}{3}*9\) so \(x=\frac{4}{3}*9=\frac{4}{\cancel{3}}*\cancel{9}^3=4*3=12\)